### v1

# Leveraging Joint Spectral and Spatial Learning with MAMBA for Multichannel Speech Enhancement

Abstract
In multichannel speech enhancement, effectively capturing spatial and spectral information across different microphones is crucial for noise reduction. Traditional methods, such as CNN or LSTM, attempt to model the temporal dynamics of full-band and sub-band spectral and spatial features. However, these approaches face limitations in fully modeling complex temporal dependencies, especially in dynamic acoustic environments. To overcome these challenges, we modify the current advanced model McNet by introducing an improved version of Mamba, a state-space model, and further propose MCMamba. MCMamba has been completely reengineered to integrate full-band and narrow-band spatial information with sub-band and full-band spectral features, providing a more comprehensive approach to modeling spatial and spectral information. Our experimental results demonstrate that MCMamba significantly improves the modeling of spatial and spectral features in multichannel speech enhancement, outperforming McNet and achieving state-of-the-art performance on the CHiME-3 dataset. Additionally, we find that Mamba performs exceptionally well in modeling spectral information.
Index Terms:
multichannel speech enhancement, spatial, spectral, state space model, Mamba
I Introduction
Multichannel speech enhancement (SE) is a critical and challenging research topic in speech signal processing. In addition to the spectral information from each channel, the spatial information captured by multiple microphone arrays is an important cue for multichannel SE. Compared to single channel methods [1, 2, 3, 4, 5, 6], multichannel techniques have the potential to significantly enhance speech quality by effectively incorporating both spatial and spectral information. Spatial information refers to the differences in the sound signal captured by multiple microphones, which helps determine the location of sound sources. This allows systems to distinguish between speech and noise based on direction. Spectral information, on the other hand, refers to the frequency content of the signal, helping separate speech from noise by analyzing different frequency bands. Combining both spatial and spectral data improves the performance of SE.
Recently, there are models specifically designed to process spatial and spectral information across multiple channels. Notably, this advancement is exemplified by mechanisms such as the channel attention approach introduced by Bahareh Tolooshams [7]. In addition, specific neural networks are designed to specifically target the spatial distinctions between narrowband speech and noise [8, 9], further optimizing spatial information processing. The spectral domain is another crucial feature for differentiating clean speech from noise. Researchers have demonstrated that both full-band and sub-band regions contain valuable spectral information [10]. The state-of-the-art performance is achieved by modeling the spatial information in narrow-band and full-band frequencies while effectively combining sub-band and full-band features [11].
Mamba [12], a recent state-space model with a selection mechanism, has demonstrated high performance in single-channel speech separation [13, 14] and SE [15]. Building on this, SpatialNet introduces a variant of Mamba tailored for long-term streaming multichannel SE [16], particularly suited for scenarios involving stationary and moving speakers. However, SpatialNet focuses primarily on utilizing spatial information. Research exploring the joint use of equally important spatial and spectral features in state-space models for multichannel SE remains unexplored.
This study proposes a novel multi-channel SE system based on Mamba, which aims to effectively capture spatial and spectral information. We develop causal and non-causal versions of the system to meet practical application requirements. The specially designed Mamba module is dedicated to modeling spatial and spectral characteristics. Extensive experiments show that on the CHiME-3 dataset, our method achieves state-of-the-art performance, outperforming McNet [11] across all SE metrics. Carefully ablation study in controlled settings reveals Mamba is particularly advantageous in spectral modeling.
II Methodology
Figure 1: (a)The basic model architecture of Mamba, (b)The proposed Uni-directional Mamba block, (c)The proposed Bi-directional Mamba block
Overview
This section first presents the Mamba framework and its two novel variants: Uni-Mamba and Bi-Mamba in Fig. 1. Subsequently, we introduce the innovative MCMamba model, which builds upon and extends the capabilities of the Mamba and Bi-Mamba architectures.
II-A Core Mamba Architecture
Figure. 1 (a) illustrates the core architecture of Mamba. Mamba integrates the H3 architecture[17] and a gated multilayer perceptron block into a stacked structure, its main components include three linear layers, a deep convolutional network, and a state-space model (SSM). Given an input X and an output Y, the system can be described by the following mathematical formulation:
$\text{Y}^{\prime}\leftarrow\text{SSM}\left(\text{SiLU}\left(\text{Conv}\left(\text{Linear}(X)\right)\right)\right)\odot\text{SiLU}\left(\text{Linear}(X)\right),$
$\text{Y}\leftarrow\text{Linear}(Y^{\prime})$
Structured State-Space Model (SSM), as in Mamba, maps an input $x$ to an output $y$ through a higher-dimensional latent state $h$, described by the following equations:
$h_{n}=\bar{A}h_{n-1}+\bar{B}x_{n},$
$y_{n}=Ch_{n}$
where $\bar{A}$ and $\bar{B}$ represent discretized state matrices. The discretization process transforms continuous parameters $(\Delta,A,B)$ into their discrete counterparts $(\bar{A},\bar{B})$, enabling the model to process discrete-time audio signals effectively.
The Structured State Space Model (SSM)[18] introduced by Mamba offers a compelling alternative to convolutional neural networks (CNNs) and recurrent neural networks (RNNs) for handling long-range dependencies in sequential data[19]. By integrating a novel input selection mechanism, the Mamba model combines the strengths of both CNNs and RNNs, significantly enhancing its capacity to model discrete data within the SSM framework. This mechanism allows the model to dynamically parameterize SSM components based on the input, selectively focusing on or ignoring specific parts of the input sequence. Consequently, Mamba excels at efficiently processing complex and extended sequences.
Figure 2: Diagram of the proposed MCMamba architecture
II-B Uni-directional and bi-directional Mamba
To address the distinct requirements of real-time and offline multichannel SE, we propose two variants of the Mamba model: Uni-Mamba in Fig.1 (b) and Bi-Mamba in Fig.1 (c). Uni-Mamba is a causal model specifically designed for real-time applications, where only the past and present information is processed. This characteristic makes Uni-Mamba particularly well-suited for real-time multichannel SE tasks, ensuring efficient performance under strict latency constraints.
A key innovation is to modify the traditional residual connection. Rather than establishing residual connections within the same dimensional space, we first augment the output dimension of the Mamba block. This expanded output is subsequently projected into a higher-dimensional space via a linear layer before forming the residual connections. This modification introduces greater flexibility and enhances the model’s capacity to capture more complex representations and dependencies within the speech data, ultimately improving the performance in multichannel SE as shown in experiments.
Bi-Mamba, in contrast, is a non-causal model designed for offline processing. Like Uni-Mamba, it employs the strategy of projecting the output of the Mamba block to a higher dimensional space before forming residual connections. The architecture utilizes two Mamba blocks: one processes the input in its original direction, while the other handles the flipped input. The outputs from both directions are combined and refined through a final linear layer. By leveraging past and future information, Bi-Mamba performs better in SE tasks.
II-C Proposed Method: MCMamba
In this work, we present MCMamba in Fig.2, a multichannel SE network specifically designed to process temporal and spectral cues through four dedicated modules: full-band spatial, narrow-band spatial, sub-band spectral, and full-band spectral. While inspired by the McNet framework, MCMamba introduces a novel variant of Mamba to further enhance its capabilities. Each module is carefully structured to progressively refine the features extracted from the preceding stage, ensuring that essential information is preserved throughout the process. MCMamba offers flexibility based on latency requirements, utilizing Bi-Mamba for non-causal (offline) processing and Uni-Mamba for causal (real-time) processing.
II-C1 spatial modeling
Spatial feature modeling in MCMamba comprises two distinct modules: the full-band spatial module (Fig. 2 (a)) and the narrow-band spatial module (Fig. 2 (b)). Each module is responsible for extracting spatial features across different frequency ranges in multichannel speech, ensuring comprehensive spatial representation across the spectrum. The spatial characteristics across the entire frequency band are captured by processing the multichannel Short-Time Fourier Transform (STFT) coefficients along the frequency axis, effectively modeling the spatial correlations across all frequencies. The input comprises both the real(Re) and imaginary(Im) components of the multichannel signal, represented by the following formula:
$x(t,f)=\begin{aligned} &[\text{Re}(X_{1}(t,f)),\text{Im}(X_{1}(t,f)),\dots,\\ &\text{Re}(X_{M}(t,f)),\text{Im}(X_{M}(t,f))]^{T}\end{aligned}$ (1)
where $M$ denotes the microphone number. Spatial cues, such as the inter-channel phase difference (IPD), are then extracted to capture the spatial relationships between the channels:
$\text{IPD}(f)=\angle X_{i}(t,f)-\angle X_{j}(t,f)$ (2)
This component does not capture temporal dependencies (always in a causal manner); therefore, the Bi-Mamba module extracts the full-band spatial features. Fully connected (FC) layers with ReLU activation are used to project the output to the desired dimension, after which the data is reshaped for further spatial processing.
The narrow-band spatial module (Fig. 2 (b)) captures the spatial information within each frequency band and models the evolution of these spatial cues. The narrow-band process below will be executed several times in parallel for each narrow band. The output from the full-band spatial module and the original noisy multichannel signal are passed to the narrow-band module in a cascade manner across the channel dimension. The time-dependent representation is formulated as follows:
$X(f)=[x(1,f),x(2,f),\dots,x(T,f)]$ (3)
By employing either the Bi-Mamba (non-causal) or Uni-Mamba (causal) modules, followed by a fully connected layer to refine spatial characteristics, MCMamba effectively divides spatial modeling into full-band and narrow-band components. This approach ensures that frequency and time domain correlations across each channel are fully leveraged, resulting in a more accurate separation of speech from noise.
II-C2 spectral modeling
Spectrum modeling in MCMamba is carried out through the sub-band and full-band modules, each designed to process the frequency and spectral information of the reference channel. The sub-band spectral module (Fig. 2 (c)) is tasked with capturing local spectral patterns and signal smoothness, both of which are critical for differentiating speech from noise. This module processes the STFT amplitude of the reference channel:
$X_{r}(t,f)=|X_{r}(t,f)|$ (4)
The data from adjacent frequencies are interconnected to utilize spectral information more effectively:
$x(t,f)=[|X_{r}(t,f-N)|,\dots,|X_{r}(t,f+N)|]$ (5)
where $N$ is set to 3 in our experiments. Depending on whether real-time or offline processing is required, either Bi-Mamba or Uni-Mamba is employed to capture the spectral dynamics over time, followed by a FC layer with ReLU activation for further refinement. The full-band spectral module (Fig. 2 (d)) integrates information across the entire frequency range to model a broad-spectrum pattern. It operates along the frequency axis, combining multiple consecutive frames to capture temporal dependencies within the broader spectral context. The model gains access to more comprehensive information by connecting these contextual frames:
$x(t,f)=[|X_{r}(t-C,f)|,\dots,|X_{r}(t,f)|]$ (6)
By utilizing both sub-band and full-band modules, MCMamba effectively captures local and global spectral cues, ensuring that spectral modeling complements spatial information to enhance the overall accuracy of speech reconstruction.
III Experiments
TABLE I: Speech Enhancement Performance Comparison. Cited from other original papers, Cited from McNet. Some papers don’t provide causal results.
Method NB-PESQ WB-PESQ STOI SDR
Performance of Non-Causal (Offline) speech enhancement
Noisy 1.82 1.27 87.0 7.5
MNMF Beamforming[20] 94.0 16.2
Oracle MVDR 2.49 1.94 97.0 17.3
CA Dense U-net[7] 2.44 18.6
Narrow-band Net[8] 2.74 2.13 95.0 16.6
FT-JNF[21] 3.17 2.48 96.2 17.7
McNet[11] 3.38 2.73 97.6 19.6
MVDR-Embedded U-Net[22] 2.64 97.4 20.6
MCMamba 3.49 2.98 98.2 20.8
Performance of Causal (Online) speech enhancement
Noisy 1.82 1.27 87.0 7.5
Narrow-band Net[8] 2.70 2.15 94.7 16.0
FT-JNF[21] 2.80 2.23 95.4 16.9
McNet[11] 3.29 2.67 97.2 19.0
MCMamba 3.41 2.89 97.6 19.8
TABLE II: Ablation study on modules for spatial and spectral modeling, where the spatial features of the narrow band in Fig. 2(b) and the spectral features of the sub-band Fig. 2(c) are time-dependent
Spatial Spectral NB-PESQ WB-PESQ STOI SDR
LSTM LSTM 3.27 2.72 97.46 19.30
Mamba LSTM 3.46 2.80 97.60 19.22
LSTM Mamba 3.52 3.02 98.10 20.50
III-A Experimental setup
Dataset We used the simulated dataset from the 3rd CHiME Challenge, comprising 7,138, 1,640, and 1,320 utterances for training, development, and testing. The data, sampled at 16 kHz using a six-microphone tablet, includes multichannel noise from four environments: bus, cafe, pedestrian area, and street intersection. The official CHiME Matlab script was used to generate the dataset. We dynamically generated samples with random noise clips and varied SNR between -5 and 10 dB to enhance training. Speech signals were simulated by adding delays to single-channel utterances without reverberation. A Hann window of 512 samples with a 256-sample step size was applied for processing.
Model configurations The output sizes of the Mamba blocks (either Uni- or Bi-) for full-band, narrow-band spatial feature, sub-band spectrum, and full-band spectrum modeling are 64, 64, 64, and 2. The corresponding hidden cell sizes are 128, 256, 384, and 128, respectively. It is worth noting that the sub-band spectrum modeling uses three neighbouring frequency bands ($N$), while the full-band spectrum modeling uses the context of $C=5$ frames.
Training setting The max epoch is 500, with the best model selected based on the validation set, with an initial learning rate of 0.001 and a decay rate of 0.992 per epoch, using the Adam optimizer. The 5th channel was selected as the reference channel, and the performance was evaluated using NB-PESQ and WB-PESQ [23], STOI [24], and SDR [25].
III-B Experimental results
III-B1 Baseline methods
In Table I, we compare our method with the following multi-channel SE techniques: MNMF Beamforming [20], which employs multi-channel non-negative matrix factorization (MNMF) followed by beamforming for SE; Oracle MVDR, which utilizes the actual noise spatial covariance matrix for enhancement; Narrow-band Net [8], which uses two LSTMs to capture narrow-band spatial information; FT-JNF [21], an improved version of Narrow-band Net that more effectively leverages full-frequency information; McNet [11] (previous SOTA), a fusion network that incorporates LSTMs to model both spatial and spectral information; and MVDR-Embedding U-Net [22], which integrates the minimum variance distortionless response (MVDR) module into the U-Net architecture, combining beamforming with neural networks for enhanced speech quality.
III-B2 Results and analysis
Table I compares offline SE results for various non-causal models. Evidently, traditional signal processing techniques, such as the MVDR-Embedded U-Net combined with deep learning methods, often yield superior outcomes. Notably, FT-JNF, which effectively leverages spatial information, outperforms Narrow-band Net. McNet, which integrates spatial and spectral features, underscores the critical importance of utilizing spectral information in multichannel SE. Our proposed MCMamba model achieves the best performance, demonstrating the strength of Uni-Mamba and Bi-Mamba in modeling spatial and spectral information across multiple channels.
Given the significance of real-time SE, we also compared the performance of causal models. MCMamba delivered highly competitive results, surpassing even all the corresponding non-causal models. This highlights the effectiveness of our Mamba module in efficiently exploiting spatial information across different bandwidths and spectral cues across frequency bands. Notably, despite the absence of future information, MCMamba achieves excellent enhancement performance, affirming the robustness of the proposed approach.
III-B3 Ablation study
An ablation study further explored Mamba’s capabilities in spatial and spectral modeling. We used LSTM, which excels at temporal modeling, to alternate with the Mamba blocks we designed for spatial and frequency modeling. Note that the models used in Table II are causal, Uni-Mamba or LSTM for involving temporal correlation. Bi-Mamba or BLSTM is used for module modeling frequency correlation. All three models in Table II are with similar parameter numbers. When spatial and spectral features are modeled using LSTM, it serves as the baseline. When Mamba is introduced for spectral or spatial modeling, all metrics improves for a large margin. Notably when Mamba is used for spectral modeling, the performance significantly improved, especially for NB-PESQ and WB-PESQ, which are highly sensitive to spectral information. This demonstrates the advantages of Mamba over LSTM in modeling spectral information, especially its more remarkable ability to capture complex spectral dynamics.
IV Conclusion
This paper proposes MCMamba, a novel model designed for multichannel speech enhancement, leveraging spatial and spectral information through Uni-Mamba and Bi-Mamba models for causal and non-causal processing. MCMamba achieves state-of-the-art in the CHiME-3 dataset. Our experiments also demonstrate that Mamba architecture significantly improves spectral modeling, outperforming traditional methods.